Evaluate Dy
2x+4y=12
3x+5y=14

Answer by stanbon(75887)

Evaluate Dy
2x+4y=12
3x+5y=14
---
Dy:
2...12
3...14
--
Dy = 2*14-12*3 = 28-36 = -8
================================
Cheers,
Stan H
==============